A culture starts with 230 bacteria. After 30 minutes the number of bacteria is 1150. 5.1 find a formula for the number of bacteria after t hours. 5.2 use the formula to find the number of bacteria after 1 hour. 5.3 after how many minutes will the number of bacteria be 125 times the original number? Leave the answer in terms of ln but in simplified form.


1
Expert's answer
2022-06-14T12:16:11-0400

The population "N(t)" at time "t" is modeled by the function 


"N(t)=N(0)e^{kt}"

5.1. Given "N(0)=230, N(0.5)=1150." Substitute


"1150=230e^{k(0.5)}"

"e^{0.5k}=5"

"k=2\ln 5"


"k=\ln 25"

"N(t)=230(25)^{t}"

5.2.


"N(1)=230(25)^{1}=5750"

5.3


"N(t)=230(25)^{t}=230(125)"

"5^{2t}=5^3"

"t=1.5"

"t=90\ minutes"